OpenStudy (anonymous):

∑∞n=1(e1/n)/n diverge or converge? ∑∞n=1(e1/n)/n diverge or converge? @Mathematics

OpenStudy (anonymous):

find using comparison test

OpenStudy (jamesj):

Can you write this out again. You're taking the sum from n = 1 to infinity of what?

OpenStudy (anonymous):

e^1/n over n

OpenStudy (jamesj):

So you want to know if \[ \sum_{i=1}^{\infty} \frac{e^{1/n}}{n} \] converges?

OpenStudy (anonymous):

yes with period from comparison test

OpenStudy (jamesj):

Right. Now does \[ \sum_1^\infty 1/n \] converge?

OpenStudy (anonymous):

no its harmonic I think

OpenStudy (jamesj):

That's right, it doesn't. Now if you can show that \[ \frac{e^{1/n}}{n} > \frac{1}{n} \] you'll have your result.